tutors can you help please
Solve 4^x-6=16^x
and
e^-2x=0.3
Thank you!!!!!!!!!!!

Answer by stanbon(48535)

Solve 4^x-6=16^x
4^(x-6) = 4^(2x)
x-6 = 2x
x = -6
--------------
and
e^-2x=0.3
Take the natural log of both sides to get:
-2x = ln0.3
-2x = -1.20397...
x = 0.60198...
cheers,
Stan H.